OpenStudy (anonymous):

What is the LCD of 5/12 and -9/16 ?

OpenStudy (anonymous):

12 have a common factor of 12,1,6,2,4and 3 16 have 1, 16,8,2,4

OpenStudy (anonymous):

thx.

OpenStudy (anonymous):

now you just need to find the least, and 1 does not work

OpenStudy (anonymous):

So basically 2

OpenStudy (anonymous):

xD

OpenStudy (anonymous):

ok, @ineedhelpnow08 can you help me with this one? Which pair of fractions is equivalent to this pair? 2/9 and 3/7 A. 2/63 and 3/63 B. 7/63 and 9/63 C. 14/63 and 27/63 D. 18/63 and 21/63

OpenStudy (anonymous):

2/9=14/63 3/7=21/63

OpenStudy (anonymous):

Please make a differnet post for each question

OpenStudy (anonymous):

all you needed to do was find the multiples to get 63 and then solve

OpenStudy (anonymous):

it be great if someone could answer my question u.u

OpenStudy (anonymous):

Send a link to it please